OpenStudy (anonymous):

Solve \[\sqrt{x-5} +7 =x\]

OpenStudy (saifoo.khan):

x=9

OpenStudy (anonymous):

i have x = -9

OpenStudy (saifoo.khan):

it can't be that. this is because suare roots can't have negative values.

OpenStudy (anonymous):

root (x - 5) + 7 = x root (x - 5) = x - 7 x - 5 = (x - 7)^2 x - 5 = x^2 - 14x + 49 0 = x^2 -15x + 54 0 = (x-9)(x-6) x = 9 and 6

OpenStudy (jamesj):

x = -9 can't be right as the term under the square root would then be negative.

OpenStudy (anonymous):

jbell and saifoo, you are missing 6.

OpenStudy (anonymous):

ohhh i see my mistake! thanks guys! :)

OpenStudy (saifoo.khan):

cow, no you are writing it wrong. LOL.

OpenStudy (anonymous):

and no i have x = 6

OpenStudy (saifoo.khan):

what you will do with \[\sqrt{1}\]??

OpenStudy (anonymous):

what \[\sqrt{1}\]?

OpenStudy (jamesj):

This is a nice problem, showing how squaring things can introduce new possibilities that don't satisfy the original equation. It's clear that x = 6 does not satisfy the equation; but it snuck in here because we squared the sqrt(x-5) term.